# Posterior Collapse as a Phase Transition in Variational Autoencoders

### v1

@@ Abstract
We investigate the phenomenon of posterior collapse in variational autoencoders (VAEs) from the perspective of statistical physics, and reveal that it constitutes a phase transition governed jointly by data structure and model hyper-parameters. By analyzing the stability of the trivial solution associated with posterior collapse, we identify a critical hyper-parameter threshold. This critical boundary, separating meaningful latent inference from collapse, is characterized by a discontinuity in the KL divergence between the approximate posterior and the prior distribution. We validate this critical behavior on both synthetic and real-world datasets, confirming the existence of a phase transition. Our results demonstrate that posterior collapse is not merely an optimization failure, but rather an emerging phase transition arising from the interplay between data structure and variational constraints. This perspective offers new insights into the trainability and representational capacity of deep generative models.
@@ 1 Introduction
Variational autoencoders (VAEs) are a class of probabilistic generative models that integrate variational inference with deep learning architectures [1, 2, 3]. Designed to learn latent representations of high-dimensional data, VAEs approximate the intractable posterior distribution over latent variables by introducing a parametric variational family, typically realized via neural networks. The learning objective is to maximize the evidence lower bound (ELBO), which balances reconstruction accuracy and the complexity of the latent encoding through a regularization term given by the Kullback–Leibler (KL) divergence between the approximate posterior and a prior distribution [1, 2, 3, 4].
VAEs have been widely and successfully applied in tasks such as image synthesis [5], anomaly detection [6], and molecular design [7]. However, as a data driven method, it is pointed out that the performance of VAEs is highly effected by data itself [8, 9, 10]. Moreover, VAEs often suffer from a phenomenon known as posterior collapse, in which a subset of the latent variables is effectively ignored [11, 12, 13, 14]. It is commonly pointed that posterior collapse reasons from the KL-divergence term in the ELBO, which encourages the approximate posterior distribution to match the prior distribution for latent variables [15, 16, 17, 18]. Several approaches have been proposed to address this issue, such as introducing a weighting coefficient $\beta$ in the ELBO ($\beta$-VAE) [19], applying anneal methods [18], or rewriting ELBO itself [20].
Recent works have provided new insights into the causes of posterior collapse by studying linear VAE models [11, 12, 13, 14]. These studies highlight alternative contributing factors, such as the the log marginal likelihood [11], the structure of the posterior distribution [12], competition among terms in the ELBO [13], and bad local extrema [14], are also responsible for posterior collapse.
In this work, we present a new perspective on posterior collapse in VAEs by framing it as a phase transition [21]. In section 3, we revisit the mathematical formulation of VAEs and derive conditions for maximizing the ELBO. Section 4 introduces our key insight: posterior collapse can be interpreted as a phase transition in VAEs, characterized by the emergence of a critical point. We validate this view using both synthetic and real-world datasets. Finally, section 5 presents a discussion and concluding remarks.
@@ 2 Related works
Linear models and analytical perspectives: To better understand the mechanism of posterior collapse, several works have studied VAEs in linear settings where analytical insights are tractable [11, 12, 13]. These studies reveal that collapse can occur even in the absence of optimization failures, due to properties intrinsic to the ELBO landscape—such as the dominance of the log marginal likelihood or the flatness of local extrema. These findings have motivated our approach, which extends beyond linear assumptions by analyzing posterior collapse as a phase transition determined jointly by data structure and model hyper-parameters.
@@ 4 Posterior collapse as phase transition
Figure 2: (a) Trained ELBO, log likelihood and posterior collapse line v.s. $\sigma^{\prime 2}$. The trained ELBO converges the posterior collapse when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}=1$. (b) Mean latent variances $\left<\sigma_{j}^{2}\right$> ($j=1,2,3,4$) v.s. $\sigma^{\prime 2}$. All of them equal 1 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}=1$. (c) KL divergence v.s $\sigma^{\prime 2}$, which converges to 0 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}=1$ (green line), and can be fitted by a quadratic function of $\sigma^{\prime 2}$ when $\sigma^{\prime 2}<\sigma^{\prime 2}_{C}=1$ (red line).
@@
has no solution other than $q_{\phi}(\bm{z}|\bm{x})=p_{\rm lat}(\bm{z})$ under some $p_{\rm data}(\bf x)$ [22]. Thus, it is reasonable to regard the onset of posterior collapse as a phase transition: from nontrivial, possible multiple extrema points to the trivial one. Another evidence for claiming posterior collapse as phase transition falls on the KL-divergence term in ELBO. The KL-divergence $D_{\rm KL}\left(q_{\phi}(\bm{z}|\bm{x})||p_{\rm lat}(\bm{z})\right)$ becomes and keeps 0 if fallen into posterior collapse, which may lead to discontinuities in the derivative of the KL divergence.
@@
where $\xi_{i}^{2}$ ($i=1,2,\cdots,N$) are the eigenvalues of the covariance matrix $\bm{\Sigma}$ for the data. We proof this criterion in the Appendix.
@@
Figure 3: Principle components of the preprocessed CIFAR10 data. The maximum one is 23.78, which corresponds to $\sigma^{\prime 2}_{C}$.
Figure 4: (a) Trained ELBO and posterior collapse line v.s. $\sigma^{\prime 2}$. The trained ELBO converges the posterior collapse when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}\approx 23.78$. (b) Mean latent variances $\left<\sigma_{j}^{2}\right$> ($j=1,\cdots,8$) v.s. $\sigma^{\prime 2}$. All of them equal 1 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}\approx 23.78$. (c) KL divergence v.s $\sigma^{\prime 2}$, which converges to 0 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}\approx 23.78$.
@@
The principle components of $\bm{x}^{\prime}$ are shown in Figure. 3, with the largest being 23.78, implying $\sigma^{\prime 2}_{C}\approx 23.78$. The architecture of the VAE model is similar to that used for the synthetic dataset, but the dimensions of hidden layers are 512 and 256 for the encoder, while 256 and 512 for the decoder. We set $n=8$ for latent variables. Similar behaviors are observed near the critical point $\sigma^{\prime 2}_{C}\approx 23.78$, as seen in the ELBO (Figure. 4(a)), the mean latent variances $\left<\sigma^{2}_{j}\right$> (Figure. 4(b)) and the KL divergence (Figure. 4(c)). We can also explore the discontinuous of KL-divergence by fitting it with the following piecewise function (red line and green line in Figure. 4(c)):
@@ 5 Conclusions
Our findings suggest that posterior collapse is not merely an artifact of optimization dynamics but a fundamental emergent behavior governed by statistical and variational constraints. This perspective not only bridges variational learning with phase transition theory but also offers practical tools for preventing posterior collapse in VAE training.
@@
Further, the role of the reconstruction noise variance $\sigma^{\prime 2}$ is analogous to the $\beta$ parameter in $\beta$-VAE, adjusting the weights of the reconstruction loss and KL-divergence in the ELBO. As such, our framework can be readily extended to $\beta$-VAEs, since introducing a $\beta$ coefficient for the KL-divergence is equivalent to rescaling the ELBO and $\sigma^{\prime 2}$, and shifting the critical point accordingly.
@@
Moreover, it is also worth exploring more precious theories to identify the posterior collapse as, and analysis the order of the phase transition, beyond discussing the stability of the trivial solutions (21)–(23). Such theories may require analytical solutions of extremes other than the trivial one, and the corresponding stability.

### v2

@@ Abstract
We investigate the phenomenon of posterior collapse in variational autoencoders (VAEs) from the perspective of statistical physics, and reveal that it constitutes a phase transition governed jointly by data structure and model hyper-parameters. By analyzing the stability of the trivial solution associated with posterior collapse, we identify a critical hyper-parameter threshold. In particular, we derive an explicit criterion for the onset of collapse: posterior collapse occurs when the decoder variance exceeds the largest eigenvalue of the data covariance matrix. This critical boundary, separating meaningful latent inference from collapse, is characterized by a discontinuity in the KL divergence between the approximate posterior and the prior distribution, where the KL divergence and its derivatives exhibit clear non-analytic behavior. We validate this critical behavior on both synthetic and real-world datasets, confirming the existence of a phase transition. The experimental results align well with our theoretical predictions, demonstrating the robustness of our collapse criterion across various VAE architectures. Our stability-based analysis demonstrate that posterior collapse is not merely an optimization failure, but rather an emerging phase transition arising from the interplay between data structure and variational constraints. This perspective offers new insights into the trainability and representational capacity of deep generative models.
@@ 1 Introduction
Variational autoencoders (VAEs) are a fundamental class of latent-variable generative models that combine variational inference with deep neural networks [1, 2, 3]. By introducing a tractable variational posterior, VAEs enable approximate Bayesian learning in high-dimensional spaces, and their training objective, evidence lower bound (ELBO), balances reconstruction fidelity with a KL-regularization term that controls the complexity of the latent representation [1, 2, 3, 4]. This framework has led to widespread success in applications including image synthesis [5], anomaly detection [6], and molecular design [7].
Despite this success, VAEs frequently exhibit posterior collapse—a phenomenon in which latent variables become uninformative and the approximate posterior degenerates to the prior [8, 9, 10, 11]. Posterior collapse poses a major limitation for representation learning, especially when the decoder is expressive or the dataset exhibits strong correlations. Classical explanations attribute collapse to the KL term in the ELBO, which encourages the posterior distribution to match the prior [12, 13, 14, 15]. Numerous heuristics such as $\beta$-VAE [16], KL annealing [15], or modified ELBO formulations [17] have been proposed, yet a principled understanding of when collapse occurs remains incomplete.
In parallel to these theoretical analyses, several recent works have explored improving the structure or expressiveness of the latent space in practical VAE applications. Dimension-weighting methods [18] enhance disentanglement by exploiting feature imbalance; Cloud-VAE [19] introduces concept-level structure into the latent representation; and PixelVAE [20] combines VAEs with autoregressive PixelCNN decoders to improve sample quality. These approaches focus on enhancing the semantic or expressive properties of the latent space, which is complementary to our goal of characterizing the conditions under which the latent variables become entirely uninformative due to posterior collapse.
Recent analytical studies have provided deeper insight by examining linear VAEs [8, 9, 10, 11]. These works reveal roles played by the log marginal likelihood [8], posterior geometry [9], competition among ELBO terms [10], and local extrema [11]. More recently, Ichikawa and Hukushima [21] analyzed a linear $\beta$-VAE using the replica method in a high-dimensional setting, deriving a phase diagram with respect to the regularization strength in ELBO and sample complexity. While these results offer valuable asymptotic characterizations, they fundamentally rely on strict linear encoder-decoder mappings and do not provide an architecture-independent condition applicable to practical non-linear VAEs.
In light of these limitations, there remains a need for a general theoretical framework that characterizes the onset of posterior collapse beyond the linear setting, and that directly connects collapse to statistical properties of the data and model hyperparameters. Our objective in this work is to provide such a framework by interpreting posterior collapse as a phase transition in the ELBO landscape.
In this paper, we make the following contributions:
  1. 1.
    We reinterpret posterior collapse as the gain of stability of the trivial ELBO extremum, providing a statistical-physics perspective on collapse in VAEs.
  2. 2.
    We derive a general and architecture-independent criterion for the onset of collapse. Specifically, for deep Gaussian VAEs, collapse occurs when the decoder variance exceeds the largest eigenvalue of the data covariance matrix.
  3. 3.
    We show that the KL divergence exhibits non-analytic behavior at the predicted critical point, consistent with phase-transition theory.
  4. 4.
    We validate the theoretical prediction using both synthetic Gaussian data and real-world datasets, demonstrating that the criterion accurately matches the empirical collapse boundary in non-linear VAEs.
The rest of this paper is organized as follows. Section 3 reviews the ELBO formulation of VAEs and characterizes the corresponding extrema. Section 4 presents the collapse criterion based on the analysis of stability, interpretes posterior collapse as a phase transition, and provides validation on synthetic and real datasets. A broader discussion of the practical implications of our findings is provided in Section 5. Section 6 concludes with discussion and future directions.
@@ 2 Related works
Linear models and analytical perspectives: To better understand the mechanism of posterior collapse, several works have studied VAEs in linear settings where analytical insights are tractable [8, 9, 10, 21]. These studies reveal that collapse can occur even in the absence of optimization failures, due to properties intrinsic to the ELBO landscape—such as the dominance of the log marginal likelihood or the flatness of local extrema. These findings have motivated our approach, which extends beyond linear assumptions by analyzing posterior collapse as a phase transition determined jointly by data structure and model hyper-parameters.
Beyond the theoretical analyses of posterior collapse, a parallel line of recent VAE research focuses on improving the structure and interpretability of the latent space. Dimension-weighting methods [18] exploit feature imbalance in VAEs to enhance disentanglement by reweighting latent dimensions, while Cloud-VAE [19] introduces concept embeddings to impose interpretable semantic structure in the latent space. Other approaches such as PixelVAE [20] combine autoregressive PixelCNN decoders with VAEs to improve expressiveness in modeling natural images. These works aim to enhance the quality, semantics, or expressiveness of latent representations. In contrast, the present study examines a different question: under what conditions does the latent representation become completely uninformative?
While prior studies have significantly advanced the understanding of posterior collapse, most analyses rely on specific structural assumptions that limit their generality. For instance, analytical treatments based on linear VAEs [8, 10, 11] offer valuable intuition but do not capture the behavior of practical non-linear models. Replica-method analyses [21] provide elegant asymptotic characterizations, yet fundamentally depend on strict linear encoder-decoder mappings and a particular high-dimensional limit. Other explanations focusing on ELBO competition or posterior mismatch [9] illuminate contributing factors but do not yield an explicit condition predicting when collapse occurs.
These limitations highlight a gap in the existing literature: there is no architecture-independent, theoretically grounded criterion that determines the onset of posterior collapse in general deep VAEs. Our work addresses this gap by analyzing the stability of the ELBO’s trivial solution and deriving a collapse condition that depends solely on the data covariance spectrum and decoder variance, independent of model architecture or linearity assumptions.
@@ 4 Posterior collapse as phase transition
Figure 2: (a) Trained ELBO, log likelihood and posterior collapse line v.s. $\sigma^{\prime 2}$. The trained ELBO converges the posterior collapse when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}=1$. (b) Mean latent variances $\left<\sigma_{j}^{2}\right$> ($j=1,2,3,4$) v.s. $\sigma^{\prime 2}$. All of them equal 1 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}=1$. (c) KL divergence v.s $\sigma^{\prime 2}$, which converges to 0 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}=1$ (green line), and can be fitted by a quadratic function of $\sigma^{\prime 2}$ when $\sigma^{\prime 2}<\sigma^{\prime 2}_{C}=1$ (red line). The fitted function is used to illustrate the non-analytic derivative behavior near the critical point.
Criterion
@@
has no solution other than $q_{\phi}(\bm{z}|\bm{x})=p_{\rm lat}(\bm{z})$ under some $p_{\rm data}(\bf x)$ [22]. Thus, it is reasonable to regard the onset of posterior collapse as a phase transition: from nontrivial, possible multiple extrema points to the trivial one. Another evidence for claiming posterior collapse as phase transition falls on the KL-divergence term in ELBO. The KL-divergence $D_{\rm KL}\left(q_{\phi}(\bm{z}|\bm{x})||p_{\rm lat}(\bm{z})\right)$ becomes and keeps 0 if fallen into posterior collapse, which may lead to discontinuities in the derivatives of the KL divergence.
@@
where $\xi_{i}^{2}$ ($i=1,2,\cdots,N$) are the eigenvalues of the covariance matrix $\bm{\Sigma}$ for the data. Intuitively, when the decoder noise dominates all principal directions of the data, the latent variables cannot encode meaningful variations, causing the trivial solution to become stable. We proof this criterion in the Appendix.
@@
It is important to note that the KL divergence is not interpreted as a thermodynamic functions, such as free energy, in our framework. Rather, it naturally appears in the ELBO and acts as an order-parameter-like quantity in the non-collapsed phase, where it is strictly positive and varies smoothly with $\sigma^{\prime 2}$. The discontinuities in its derivatives at the critical point $\sigma^{\prime 2}_{C}$ signal the transition to the collapsed phase, where the KL divergence becomes identically zero. Thus, it serves as a useful indicator for identifying and characterizing the phase transition associated with posterior collapse.
(a) Preprocessed CIFAR10
(b) Preprocessed Fashion–MNIST
Figure 3: (a) Principle components of the preprocessed CIFAR10 data. The maximum one is 23.78, which corresponds to $\sigma^{\prime 2}_{C}$. (b) Principle components of the preprocessed Fashion-MNIST data. The maximum one is 22.15.
Validation on real-world dataset
@@
The principle components of $\bm{x}^{\prime}$ are shown in Figure. 3(a), with the largest being 23.78, implying $\sigma^{\prime 2}_{C}\approx 23.78$. The architecture of the VAE model is similar to that used for the synthetic dataset, but the dimensions of hidden layers are 512 and 256 for the encoder, while 256 and 512 for the decoder. We set $n=8$ for latent variables. Similar behaviors are observed near the critical point $\sigma^{\prime 2}_{C}\approx 23.78$, as seen in the ELBO (Figure. 4(a)), the mean latent variances $\left<\sigma^{2}_{j}\right$> (Figure. 4(b)) and the KL divergence (Figure. 4(c)). We can also explore the discontinuous of KL-divergence by fitting it with the following piecewise function (red line and green line in Figure. 4(c)):
@@
Figure 4: (a) Trained ELBO and posterior collapse line v.s. $\sigma^{\prime 2}$. The trained ELBO converges the posterior collapse when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}\approx 23.78$. (b) Mean latent variances $\left<\sigma_{j}^{2}\right$> ($j=1,\cdots,8$) v.s. $\sigma^{\prime 2}$. All of them equal 1 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}\approx 23.78$. (c) KL divergence v.s $\sigma^{\prime 2}$, which converges to 0 when $\sigma^{\prime 2}>\sigma^{\prime 2}_{C}\approx 23.78$. The fitted function is used to illustrate the non-analytic derivative behavior near the critical point.
Validation of $\beta$-VAE and architecture independence
$\beta$-VAE [16] introduces a hyper-parameter $\beta$ to balance the reconstruction loss and KL-divergence in the ELBO:
$\mathcal{L}_{\beta\text{-VAE}}$ $\equiv\mathbb{E}_{p_{\rm data}(\bm{x})}\left[\mathbb{E}_{\bm{z}\sim q_{\phi}(\bm{z}|\bm{x})}\left[\ln p_{\theta}(\bm{x}|\bm{z})\right]\right]-\beta\mathbb{E}_{p_{\rm data}(\bm{x})}\left[D_{\rm KL}\left(q_{\phi}(\bm{z}|\bm{x})||p_{\rm lat}(\bm{z})\right)\right].$ (32)
By substituting Eqs. (14)–(16) into Eq. (32), we have the training target for deep Gaussian $\beta$-VAEs:
$\mathcal{L}_{\beta\text{-VAE}}^{\rm G}$ $\equiv N\ln\frac{1}{\sqrt{2\pi}\sigma^{\prime}}-\int\frac{1}{2\sigma^{\prime 2}}\sum_{i=1}^{N}\left[x_{i}-f_{i}(\bm{z})\right]^{2}q_{\phi}(\bm{z}|\bm{x})p_{\rm data}(\bm{x}){\rm d}\bm{x}{\rm d}\bm{z}$
$\qquad-\frac{\beta}{2}\int\sum_{j=1}^{n}\left[\sigma^{2}_{j}(\bm{x})+\mu_{j}^{2}(\bm{x})-1-\ln\sigma_{j}^{2}(\bm{x})\right]p_{\rm data}(\bm{x}){\rm d}\bm{x},$ (33)
which takes a similar structure as Eq. (17). Indeed, with the replacement
$\sigma^{\prime 2}\rightarrow\beta\sigma^{\prime 2},$ (34)
$\mathcal{L}_{\beta\text{-VAE}}^{\rm G}$ can be regarded as $\mathcal{L}_{VAE}^{\rm G}$ for vanilla deep Gaussian VAEs scaling by a factor $\beta$. This relationship is also implied in previous analysis of linear VAEs [9, 10]. Thus, the criterion for posterior collapse in $\beta$-VAE reads:
$\beta\sigma^{\prime 2}>\max\left[\xi_{1}^{2},\cdots,\xi_{N}^{2}\right].$ (35)
We validate the criterion (35) using Fashion-MNIST [27], which consists of 70,000 grayscale images of size $28\times 28$ ($N=784$). We preprocess the dataset with the same method in CIFAR10. The principle components are shown in Figure. 3(b), with the largest being 22.15.
To demonstrate that our criterion, Eq. (26) or (35), does not depend on the specific neural architecture, we employ convolutional encoder and decoder networks in this $\beta$-VAE. The encoder includes two convolutional layers with ReLU activations, while the decoder consists of two transposed convolutional layers, also activated with ReLU. We set $n=8$ for latent variables and $\sigma^{\prime 2}=1$, making the critical point $\beta_{C}\approx 22.15$. Similar behaviors are observed near the critical point $\beta_{C}\approx 22.15$, as seen in the ELBO (Figure. 5(a)), the mean latent variances $\left<\sigma^{2}_{j}\right$> (Figure. 5(b)) and the KL divergence (Figure. 5(c)). We can also explore the discontinuous of KL-divergence by fitting it with the following piecewise function (red line and green line in Figure. 5(c)):
$f(\beta)=\begin{cases}4.65e^{-0.18\beta}-0.09&\left(\beta<\beta_{C}\approx 22.15\right)\\ 0&\left(\beta>\beta_{C}\approx 22.15\right)\end{cases},$ (36)
of which the derivatives also show discontinuities at $\beta=\beta_{C}$.
Figure 5: (a) Trained ELBO and posterior collapse line v.s. $\beta$. The trained ELBO converges the posterior collapse when $\beta>\beta_{C}\approx 22.15$. (b) Mean latent variances $\left<\sigma_{j}^{2}\right$> ($j=1,\cdots,8$) v.s. $\beta$. All of them equal 1 when $\beta>\beta_{C}\approx 22.15$. (c) KL divergence v.s $\beta$, which converges to 0 when $\beta>\beta_{C}\approx 22.15$. The fitted function is used to illustrate the non-analytic derivative behavior near the critical point.
5 Applicability to real-world problems
Although the analysis developed in this work is rooted in the mathematical structure of the ELBO, the resulting collapse criterion has several direct implications for practical machine learning tasks involving VAEs.
Hyperparameter selection and model design
The collapse boundary (26) provides a principled guideline for avoiding posterior collapse. In practice, this means that a practitioner may inspect the covariance spectrum of a dataset and select decoder noise levels or KL regularization strengths that ensure the stability of the informative solution. This offers a diagnostic tool for determining when collapse is likely to appear during training, without the need for extensive trial-and-error hyperparameter search.
Applicability to non-linear and convolutional architectures
Because our stability analysis depends only on the local behavior of the ELBO around the trivial extremum, the resulting criterion is independent of the specific parameterization of the encoder and decoder networks. This includes widely used non-linear architectures such as convolutional VAEs in image modeling and anomaly detection. Our empirical results confirm that the predicted collapse threshold closely matches the observed behavior even when the encoder and decoder are highly expressive neural networks.
Relevance to practical domains
Posterior collapse is a common bottleneck in real-world generative modeling tasks, including representation learning, semi-supervised learning, image synthesis, and industrial anomaly detection. The stability-based framework presented here offers a quantitative interpretation of when these failures occur, and provides insight into how the architecture or hyperparameters may be adjusted to maintain informative latent representations.
@@ 6 Conclusions
Our findings suggest that posterior collapse is not merely an artifact of optimization dynamics but a fundamental emergent behavior governed by statistical and variational constraints. This perspective not only bridges variational learning with phase transition theory but also offers practical tools for preventing posterior collapse in VAE training. The KL divergence therefore provides a convenient empirical indicator of the transition: although it becomes exactly zero in the collapsed phase, the one-sided derivatives approaching the critical point from the non-trivial phase encode the non-analytic structure characteristic of phase transitions.
@@
As we have mentioned in Sec. 4, the role of the reconstruction noise variance $\sigma^{\prime 2}$ is analogous to the $\beta$ parameter in $\beta$-VAE, adjusting the weights of the reconstruction loss and KL-divergence in the ELBO. Thus, introducing a $\beta$ coefficient for the KL-divergence is equivalent to rescaling the ELBO and $\sigma^{\prime 2}$, and shifting the critical point accordingly.
@@
Moreover, it is also worth exploring more precious theories to identify the posterior collapse as well as analysis the order of the phase transition, beyond discussing the stability of the trivial solutions (21)–(23). Such theories may require analytical solutions of extremes other than the trivial one, and the corresponding stability.